January 29th 2024

From Hall A Wiki
Jump to: navigation, search
  • Updates
    • HCal
      • looked at cabling fror HCAL trigger
      • Question LED calibration : LED was removed from ePAS -
      • HCal trigger firmware : Ben starting merging current VTP readout firware with HCal trigger
      • RF timing - check before start of GEn RP and with beam
    • CDet
      • Ralph and Angelo working on last half module running and taking cosmics - will be able to get resistors values to equalize the gain
      • installation of level translator in spectrometer -vertical rail still need to be mounted - Zac will help to set it up
      • will make the VETROC adapters to ECL - will check with electronics group
      • will make long ECL cables
    • GEn RP
      • David will contact Alex
      • V1190 for active analyzer will be sent from Glasgow
      • need to check with Bill for status of FADC and cabling
  • Digital trigger for GEn RP
    • BBCal : Arun willing to help for BBCAL FADC trigger
    • implement coincidence HCAL and BBCAL in BBCAL VTP
    • need to send BBCAL trigger and HCAL trigger to TS for cross checks


  • Software review GEp